8) A basic cell phone plan at Verizon Wireless costs $45 a month with an additional $0.10 per text message sent. AT&T charges $30 a month with an additional $0.13 per text message sent. How many text messages would you have to send for the plans to cost the same? If you send an average of 1200 text messages a month, which company would you want to choose as your provider?
a. Define your variables
(I know that X=# of Verizon texts, Y=# of AT&T texts)
b. Write your system of equations (X<=0,Y<=0,?,?)
c. Show your work
d. Answer the question in a complete sentence.
------------
I'm not sure how to write out the system of equations? Do I exclude the amount they charge per month? Is it .10x<=1200 and .13y<=1200?

Answer by mananth(16949)


Number of tests made =x
Verzon C(x) = 45 + 0.1x => 45+0.1*1200 =165

AT & T C(y) = 30 + 0.13x=> 30 + 0.13*1200=186
For 1200 texts C(x) is less than C(y)
If the costs are to be the same then

45 + 0.1x = 30 + 0.13x
15=0.03x
15/0.03= x
500=x
500 texts